hey there i need help i mean SERIOUS help on graping systems of equation...here is the equation {y=x+3
y=x-1
i need to graph such thing like that and i dnt got it plz help

Answer by Edwin McCravy(20055)

I'll just graph one line:
y=x+3
Make up a number, any number. It can be any number if your graph
is big enough, but make is small so you won't need a big graph.
Let's pick a number to substitute for x. say 2. Substitute 2 for
x in
y = x + 3
y = 2 + 3
y = 5
So that means when x is 2, y is 5. So that makes a point (2,5).
Plot that point. It's in line with 2 on the x-axis and it's also in
line with 5 on the y axis.


Make up aother number, any number. Let's pick -1. Substitute -1 for
x in
y = x + 3
y = -1 + 3
y = 2
So that means when x is -1, y is 2. So that makes a point (-1,2).
Plot that point. It's in line with -1 on the x-axis and it's also in
line with 2 on the y axis.

Make up aother number, any number. Let's pick 0. Substitute 0 for
x in
y = x + 3
y = 0 + 3
y = 3
So that means when x is 0, y is 3. So that makes a point (0,3).
Plot that point. Now since x is 0, that ,eans that the point is ON
the y-axis at the number 3, like this.

Make up aother number, any number. Let's pick -5 this time.
Substitute -5 for x in
y = x + 3
y = -5 + 3
y = -2
So that means when x is -5, y is -2. So that makes a point (-5,-2).
Plot that point. It's in line with -5 on the x-axis and it's also in
line with -2 on the y axis.

)}}}
We could keep on getting points, but that'e nough. Now take a
ruler and draw a line through those points. They will always
line up line that if you don't make a mistake.

Now do the same thing with the other equation,
y=x-1
Plot some points on the same graph on which you drew
the other one, and you'll gat another line parallel to that one.
Edwin